# Square Functions Controlling Smoothness and Higher-Order Quantitative Rectifiability

### v1

@@ 1. Introduction
(1.1) $\beta_{2,k}(X,r)=\inf_{\text{$P$ a $k$-plane}}\left(\fint_{B_{r}(X)}\left(\frac{\operatorname{dist}(Y,P)}{r}\right)^{2}\,\mathrm{d}\mu(y)\right)^{\frac{1}{2}}$
@@
(1.2) $\int_{0}^{R}\int_{B_{R}(X)\cap\Sigma}\beta_{2,k}(Y,r)\frac{\,\mathrm{d}\mu(Y)\,\mathrm{d}r}{r}\leq CR^{k},\,\,\,\,\,\,\,\,\,\,\text{for all$X\in\Sigma$, $R\in(0,\operatorname{diam}\Sigma)$}$
@@
$\int_{0}^{1}\beta_{2,k}(X,r)^{2}\frac{\,\mathrm{d}r}{r}<\infty\,\,\,\,\text{$\mu$-a.e. $X\in\operatorname{supp}(\mu)$}$
@@
$\int_{0}^{\infty}\beta_{2,k}(X,r)^{2}\frac{\,\mathrm{d}r}{r}<\infty\,\,\,\,\text{ for $\mu$-a.e. $X\in\mathbb{R}^{d}$}$
@@
$\int_{0}^{1}\beta_{2,k}(X,r)^{2}\frac{\,\mathrm{d}r}{r}<\infty\,\,\,\,\text{$\mu$-a.e. $X\in\operatorname{supp}(\mu)$}$
@@
(1.7) $\int_{0}^{1}\left(\frac{\beta_{2,k}(X,r)}{r^{\alpha}}\right)^{2}\frac{\,\mathrm{d}r}{r}<\infty$
@@
(1.9) $\int_{0}^{1}\left(\frac{\beta_{2,\infty}(X,r)}{r^{\alpha}}\right)^{2}\frac{\,\mathrm{d}r}{r}<\infty$
@@
(1.11) $\limsup_{r\rightarrow\infty}r^{-\alpha}\beta_{p}(X,r)<\infty$
@@
$\nu_{0}^{f}(x,r)=\inf_{\text{$c$ a constant}}\left(\fint_{B_{r}(x)}\left(\frac{|f(y)-c|}{r}\right)^{2}\,\mathrm{d}y\right)^{\frac{1}{2}}$
@@
$\nu_{1}^{f}(x,r)=\inf_{\text{$l$ linear}}\left(\fint_{B_{r}(x)}\left(\frac{|f(y)-l(y)|}{r}\right)^{2}\,\mathrm{d}y\right)^{\frac{1}{2}}$
@@
(1.13) $\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\nu_{0}^{f}(y,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}y\,\mathrm{d}r}{r}\leq CR^{d}$
@@
(1.14) $\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\nu_{1}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\leq CR^{d}$
@@
(1.15) $\int\frac{|f(x)|}{1+|x|^{d+\epsilon}}\,\mathrm{d}x$
@@
(1.17) $\int_{0}^{R}\int_{\Delta_{R}(Z)}\left(\frac{\beta_{2,d}(X,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}\mathcal{H}^{d}|_{\Gamma}(x)\,\mathrm{d}r}{r}\leq CR^{d}$
@@
BMO-Sobolev spaces play a central role in the theory of parabolic geometric measure theory and boundary value problems for the heat equation in rough domains. Let us refer to Lip$(1,1/2)$ graphs (i.e. graphs given by functions which are Lipschitz with respect to the parabolic norm) parabolic Lipschitz graphs, and Lip$(1,1/2)$ graphs with underlying $(1/2)$-time derivative in parabolic BMO regular parabolic Lipschitz graphs. We refer the reader to a recent paper of Dindoš and Sätterqvist [DSa] where the authors show that the space of such functions can be characterized through a parabolic version of the BMO-Sobolev norm developed by Strichartz in [Str]. It was long conjectured, and recently proved by Bortz, Hofmann, Martell, and Nyström in [BHMN], that solvability of the $L^{p}$ Dirichlet problem for the heat equation in a parabolic Lipschitz graph domain implies that the domain is a regular parabolic Lipschitz graph domain. The converse to this statement was proved in 1995 by Lewis and Murray [LM], where they showed that the $L^{p}$ Dirichlet problem for the heat equation is necessarily solvable for some $p$ (which is possibly very large) in regular parabolic Lipschitz graph domains. This result was further refined by Hofmann and Lewis in 1996 [HL], where they showed that, if the underlying function of a regular parabolic Lipschitz graph domain has sufficiently small $BMO$ norm of the $1/2$ time derivative, then the $L^{2}$ Dirichlet problem for the heat equation is solvable. Moreover, in [HL] the authors show that the aforementioned result is sharp by producing a regular parabolic Lipschitz graph domain for which the $L^{2}$ Dirichlet problem for the heat equation is not solvable. The same authors prove analogous results for the Regularty and Neumann problems in regular parabolic Lipschitz graph domains in [HL2] and [HL3]. In [DPP], Dindoš, Petermichl, and Pipher show two results. First, in regular parabolic Lipschitz domains, scale-invariant mutual absolute continuity of certain parabolic measures with respect to surface measure implies BMO-solvability of the parabolic PDE associated to the aforementioned parabolic measure, and in addition the modulus of the gradient of solutions which are continuous on the boundary must be controlled near the boundary, in a Carleson measure sense. They also prove that similar Carleson measure control on gradients of solutions which are continuous on the boundary implies $L^{p}$ solvabillity of the Dirchlet problem for the parabolic equations they consider. In [DDH], Dindoš, Dyer, and Hwang prove $L^{p}$ solvability of certain parabolic equations in regular parabolic Lipschitz graph domains. The condition that the half-time derivative be in parabolic $BMO$ is also a sufficient condition for boundedness of a certain singular integral, see [H2].
The notion of parabolic uniform rectifiability was introduced by Hofmann, Lewis, and Nyström in [HLN], where they prove results about “big pieces” approximation by regular parabolic Lipschitz graphs for Reifenberg flat sets satisfying a parabolic version of the Carleson measure condition (1.2). In [HLN], the same authors use the aforementioned result to establish mutual absolute continuity of the surface measure and caloric measure in such domains. In recent years there has been an increased interest in the theory of parabolic uniform rectifiability. In [GH], Genschaw and Hofmann show that a local ampless condition for certain parabolic measures implies a weak reverse Hölder inequality on the Poisson Kernel in domains with boundary satisfying certain time-backward thickness conditions. In [BHHLN], Bortz, the author of this paper, Hofmann, Luna-García, and Nyström prove a parabolic analog of a result of David and Jerison [DJ], which shows, among other things, that certain parabolic uniformly rectifiable domains satisfying a two-sided corkscrew condition and a time backwards and forwards thickness condition necessarily have big pieces of regular parabolic Lipschitz graphs. This result, together with the result of [GH] implies that the $L^{p}$ Dirichlet problem for the heat equation is solvable in such domains, at least for sufficiently large $p$. In [BHHLN2], the authors show that domains with parabolic uniformly rectifiable boundaries satisfy “Carleson measure estimates,” in the sense that solutions to the heat equation in these domains have gradients which are not too large near the boundary in a Carleson measure sense. In [BHHLN3], the authors prove that parabolic uniformly rectifiable sets admit “corona decompositions” in terms of regular parabolic Lipschitz graphs. Such decompositions were established in the “Euclidean” case by David and Semmes [DS1], and have served as a crucial tool for proving the equivalence of many different conditions to uniform rectifiability.
@@ 2. Notation, Definitions
Of particular interest to us is the space $BMO$
@@
$\|f\|_{*}:=\sup_{x\in\mathbb{R}^{d},r>0}\fint_{B_{r}(x)}|f(y)-\fint_{B_{r}(x)}f(z)\,\mathrm{d}z|\,\mathrm{d}y<\infty$
@@
$\langle f\rangle_{B_{r}(x)}=\fint_{B_{r}(z)}f(z)\,\mathrm{d}z$
@@
$I_{\alpha}(\varphi)(x)=c_{d,\alpha}p.v.\int_{\mathbb{R}^{d}}\frac{\varphi(y)}{|x-y|^{d-\alpha}}\,\mathrm{d}y$
@@
$I_{\alpha}(\varphi)(x)=\big{(}(2\pi|\xi|)^{-\alpha}\hat{\varphi}(\xi)\check{\big{)}}(x)$
@@
$(D_{\alpha}\varphi)(x):=(2\pi)^{\alpha}(|\xi|^{\alpha}\hat{\varphi}(\xi)\check{)}(x)$
@@
$\int x^{\gamma}\varphi(x)\,\mathrm{d}x=0$
@@
We are now ready to define the space $I_{\alpha}(BMO)$
@@
$I_{\alpha}(b)(\varphi):=\int b(x)I_{\alpha}(\varphi)(x)\,\mathrm{d}x$
@@
For $10<\alpha<$ 1, (equivalence classes of) functions in $I_{\alpha}(BMO)$ are Hölder continuous with exponent $\alpha$. For $21<\alpha<$ 2, (equivalence classes of) functions in $I_{\alpha}(BMO)$ are differentiable with $(\alpha-1)$-Hölder continuous gradients. If $\alpha=1$, (equivalence classes of) functions in $I_{\alpha}(BMO)$ are differentiable almost everywhere with gradients in $BMO$.
$[f]_{\alpha}=\sup_{x\neq y}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}$
@@ 3. $I_{\alpha}(BMO)$ Implies Square Function Bounds
In this section we show that functions in $I_{\alpha}(BMO)$ necessarily satisfy the square function bounds (1.13) and (1.14) for $10<\alpha<$ 1 and $21\leq\alpha<$ 2, respectively. We begin by showing that localizations of $I_{\alpha}(BMO)$ functions satisfy appropriate bounds which comport with the scale of the localizations. As we shall see, the case when $\alpha=1$ is fundamentally different. We should note that Propositions 3.1 and 3.4 are somewhat similar to a Lemma proved in the appendix of [H2], although our case is somewhat more difficult because we need to analyze the singular integral which makes an appearance in the final term of our estimate in the case that $21\leq\alpha<$ 2.
@@
$g(x):=\phi\left(\frac{x-z}{R}\right)\big{(}f(x)-f(z)-\nabla f(z)\cdot(x-z)\big{)}$
@@
(3.2) $\int_{\mathbb{R}^{d}}\big{|}|\xi|^{\alpha}\hat{g}(\xi)\big{|}^{2}\,\mathrm{d}\xi\leq CR^{d}$
@@
$\int_{\mathbb{R}^{d}}\left(\hat{g}(\xi)|\xi|^{\alpha}\right)^{2}\,\mathrm{d}\xi\approx\int_{\mathbb{R}^{d}}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}\xi$
[…]
$\int_{|x-z|\geq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x+\int_{|x-z|<100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x=I+II$
@@
(3.3) $I=\int_{|x-z|\geq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ =\int_{|x-z|\geq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{\phi\left(\frac{y-z}{R}\right)\big{(}f(y)-f(z)-\nabla f(z)\cdot(y-z)\big{)}}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ \lesssim_{[\nabla f]_{\alpha-1}^{2}}R^{2\alpha}\int_{|x-z|\geq 100R}\left(\int_{|y-z|\leq 4R}\frac{1}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ \lesssim R^{2d+2\alpha}\int_{|x-z|\geq 100R}\frac{1}{|x-z|^{2d+2\alpha}}\,\mathrm{d}x\approx R^{d}$
@@
$g(x)-g(y)=\left(\phi\left(\frac{x-z}{R}\right)-\phi\left(\frac{y-z}{R}\right)\right)\left(f(x)-f(z)-\nabla f(z)\cdot(x-z)\right)\\ -\phi\left(\frac{y-z}{R}\right)\left(f(y)-f(x)-\nabla f(z)\cdot(y-x)\right)$
@@
$II\lesssim\int_{|x-z|<100R}\left(p.v.\int_{|x-y|\leq 10R}\frac{f(x)-f(z)-\nabla f(z)\cdot(x-z)}{|x-y|^{d+\alpha}}\left(\phi\left(\frac{x-z}{R}\right)-\phi\left(\frac{y-z}{R}\right)\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ +\int_{|x-z|<100R}\left(\int_{|x-y|\geq 10R}\frac{f(x)-f(z)-\nabla f(z)\cdot(x-z)}{|x-y|^{d+\alpha}}\left(\phi\left(\frac{x-z}{R}\right)-\phi\left(\frac{y-z}{R}\right)\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ +\int_{|x-z|<100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{f(y)-f(x)-\nabla f(z)\cdot(y-x)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ =:III+IV+V$
@@
$|f(x)-f(z)-\nabla f(z)\cdot(x-z)|\lesssim_{[\nabla f]_{\alpha-1}}R^{\alpha}$
$\left|\phi\left(\frac{y-z}{R}\right)-\phi\left(\frac{x-z}{R}\right)-\nabla\left(\phi\left(\frac{x-z}{R}\right)\right)\cdot(y-x)\right|\lesssim\frac{|x-y|^{2}}{R^{2}}$
@@
$\int_{|x-z|<100R}\bigg{(}p.v.\int_{|x-y|\leq 10R}\frac{f(x)-f(z)-\nabla f(z)\cdot(x-z)}{|x-y|^{d+\alpha}}\\ \cdot\left(\phi\left(\frac{x-z}{R}\right)+\nabla\left(\phi\left(\frac{x-z}{R}\right)\right)\cdot(y-x)-\phi\left(\frac{y-z}{R}\right)\right)\,\mathrm{d}y\bigg{)}^{2}\,\mathrm{d}x$
@@
$[\nabla f]_{\alpha-1}^{2}\int_{|x-z|<100R}R^{2\alpha-4}\left(\int_{|x-y|\leq 10R}\frac{1}{|x-y|^{d+\alpha-2}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\lesssim[\nabla f]_{\alpha-1}^{2}R^{d}$
@@
$\left|\phi\left(\frac{x-z}{R}\right)-\phi\left(\frac{y-z}{R}\right)\right|\lesssim\frac{|x-y|}{R}$
@@
$IV\lesssim_{[\nabla f]_{\alpha-1}^{2}}\int_{|x-z|<100R}R^{2\alpha-2}\left(\int_{|x-y|\geq 10R}\frac{1}{|x-y|^{d+\alpha-1}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\lesssim R^{d}$
@@
$V\lesssim\int_{|x-z|<100R}\left(\int_{|x-y|\geq 10R}\frac{(\nabla f(z)-\nabla f(x))\cdot(x-y)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ +\int_{|x-z|<100R}\left(p.v.\int_{|x-y|\leq 10R}\frac{(\nabla f(z)-\nabla f(x))\cdot(x-y)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ +\int_{|x-z|<100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{f(y)-f(x)-\nabla f(x)\cdot(y-x)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ =:VI+VII+VIII$

### v2

@@ 1. Introduction
(1.1) $\beta_{2,k}(X,r)=\inf_{\text{$P$ a $k$-plane}}\left(\fint_{B_{r}(X)}\left(\frac{\operatorname{dist}(Y,P)}{r}\right)^{2}\,\mathrm{d}\mu(y)\right)^{\frac{1}{2}}.$
@@
(1.2) $\int_{0}^{R}\int_{B_{R}(X)\cap\Sigma}\beta_{2,k}(Y,r)^{2}\frac{\,\mathrm{d}\mu(Y)\,\mathrm{d}r}{r}\leq CR^{k},\,\,\,\,\,\,\,\,\,\,\text{for all$X\in\Sigma$, $R\in(0,\operatorname{diam}\Sigma)$}.$
We also define, for $k\in\{1,...,d-1\}$ the upper $k$-density and lower $k$-density, respectively, of a measure $\mu$ at a point $X\in\operatorname{supp}(\mu)$
$\theta^{k,*}(\mu)(X)=\limsup_{r\rightarrow 0^{+}}\frac{\mu(B_{r}(X))}{(2r)^{k}},$
$\theta^{k}_{*}(\mu)(X)=\liminf_{r\rightarrow 0^{+}}\frac{\mu(B_{r}(X))}{(2r)^{k}}.$
For $k\in\{1,...,d-1\}$, a say that a measure $\mu$ is $k$-rectifiable if there are Lipschitz functions $f_{i}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{d}$ such that
$\mu\bigg{(}\mathbb{R}^{d}\setminus\bigcup_{i}f_{i}(\mathbb{R}^{n})\bigg{)}=0.$
The definition of $C^{1,\alpha}$ rectifiability is nearly identical, the only difference being that the Lipschitz functions $f_{i}$ are replaced by $C^{1,\alpha}$ functions.
@@
$\int_{0}^{1}\beta_{2,k}(X,r)^{2}\frac{\,\mathrm{d}r}{r}<\infty\,\,\,\,\text{$\mu$-a.e. $X\in\operatorname{supp}(\mu)$},$
@@
$\int_{0}^{\infty}\beta_{2,k}(X,r)^{2}\frac{\,\mathrm{d}r}{r}<\infty\,\,\,\,\text{ for $\mu$-a.e. $X\in\mathbb{R}^{d}$}.$
@@
$\int_{0}^{1}\beta_{2,k}(X,r)^{2}\frac{\,\mathrm{d}r}{r}<\infty\,\,\,\,\text{$\mu$-a.e. $X\in\operatorname{supp}(\mu)$},$
@@
(1.7) $\int_{0}^{1}\left(\frac{\beta_{2,k}(X,r)}{r^{\alpha}}\right)^{2}\frac{\,\mathrm{d}r}{r}<\infty.$
@@
(1.9) $\int_{0}^{1}\left(\frac{\beta_{2,\infty}(X,r)}{r^{\alpha}}\right)^{2}\frac{\,\mathrm{d}r}{r}<\infty.$
@@
(1.11) $\limsup_{r\rightarrow\infty}r^{-\alpha}\beta_{p}(X,r)<\infty.$
@@
$\nu_{0}^{f}(x,r)=\inf_{\text{$c$ a constant}}\left(\fint_{B_{r}(x)}\left(\frac{|f(y)-c|}{r}\right)^{2}\,\mathrm{d}y\right)^{\frac{1}{2}},$
@@
$\nu_{1}^{f}(x,r)=\inf_{\text{$l$ affine}}\left(\fint_{B_{r}(x)}\left(\frac{|f(y)-l(y)|}{r}\right)^{2}\,\mathrm{d}y\right)^{\frac{1}{2}}.$
@@
(1.13) $\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\nu_{0}^{f}(y,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}y\,\mathrm{d}r}{r}\leq CR^{d}.$
@@
(1.14) $\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\nu_{1}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\leq CR^{d},$
@@
(1.15) $\int\frac{|f(x)|}{1+|x|^{d+\epsilon}}\,\mathrm{d}x.$
@@
(1.17) $\int_{0}^{R}\int_{\Delta_{R}(Z)}\left(\frac{\beta_{2,d}(X,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}\mathcal{H}^{d}|_{\Gamma}(X)\,\mathrm{d}r}{r}\leq CR^{d}.$
@@
BMO-Sobolev spaces play a central role in the theory of parabolic uniform rectifiability. Let us refer to Lip$(1,1/2)$ graphs (i.e. graphs given by functions which are Lipschitz with respect to the parabolic norm) parabolic Lipschitz graphs, and Lip$(1,1/2)$ graphs with underlying $(1/2)$-time derivative in parabolic BMO regular parabolic Lipschitz graphs. We refer the reader to a recent paper of Dindoš and Sätterqvist [DSa] where the authors show that regular parabolic Lipschitz functions can be characterized through a parabolic version of the BMO-Sobolev norm developed by Strichartz in [Str]. In the theory of parabolic uniform rectifiability, parabolic regular Lipschitz graphs play a role analogous to that of Lipschitz graphs in the standard theory of uniform rectifiability.
@@ 2. Notation, Definitions
Of particular interest to us is the space $BMO$.
@@
$\|f\|_{*}:=\sup_{x\in\mathbb{R}^{d},r>0}\fint_{B_{r}(x)}|f(y)-\fint_{B_{r}(x)}f(z)\,\mathrm{d}z|\,\mathrm{d}y<\infty.$
@@
$\langle f\rangle_{B_{r}(x)}=\fint_{B_{r}(z)}f(z)\,\mathrm{d}z.$
@@
$I_{\alpha}(\varphi)(x)=c_{d,\alpha}\int_{\mathbb{R}^{d}}\frac{\varphi(y)}{|x-y|^{d-\alpha}}\,\mathrm{d}y,$
@@
$I_{\alpha}(\varphi)(x)=\big{(}(2\pi|\xi|)^{-\alpha}\hat{\varphi}(\xi)\check{\big{)}}(x),$
@@
$(D_{\alpha}\varphi)(x):=(2\pi)^{\alpha}(|\xi|^{\alpha}\hat{\varphi}(\xi)\check{)}(x).$
@@
$\int x^{\gamma}\varphi(x)\,\mathrm{d}x=0.$
@@
We are now ready to define the space $I_{\alpha}(BMO)$.
@@
$I_{\alpha}(b)(\varphi):=\int b(x)I_{\alpha}(\varphi)(x)\,\mathrm{d}x.$
@@
For $10<\alpha<$ 1, (equivalence classes of) functions in $I_{\alpha}(BMO)$ are Hölder continuous with exponent $\alpha$. For $21<\alpha<$ 2, (equivalence classes of) functions in $I_{\alpha}(BMO)$ are differentiable with $(\alpha-1)$-Hölder continuous gradients. If $\alpha=1$, (equivalence classes of) functions in $I_{\alpha}(BMO)$ are differentiable almost everywhere with gradients in $BMO$. In particular, $[f]_{\alpha}\lesssim\|D_{\alpha}f\|_{*}$ for $\alpha\in(0,1)$, and $[\nabla f]_{\alpha-1}\lesssim\|D_{\alpha}f\|_{*}$ for $21<\alpha<$ 2.
@@ 3. $I_{\alpha}(BMO)$ Implies Square Function Bounds
In this section we show that functions in $I_{\alpha}(BMO)$ necessarily satisfy the square function bounds (1.13) and (1.14) for $10<\alpha<$ 1 and $21\leq\alpha<$ 2, respectively. We begin by showing that localizations of $I_{\alpha}(BMO)$ functions satisfy appropriate bounds which comport with the scale of the localizations. As we shall see, the case when $\alpha=1$ is fundamentally different. We should note that Propositions 3.1 and 3.5 are somewhat similar to a Lemma proved in the appendix of [H2], although our case is somewhat more difficult because we need to analyze the singular integral which makes an appearance in the final term of our estimate in the case that $21\leq\alpha<$ 2.
@@
$g(x):=\phi\left(\frac{x-z}{R}\right)\big{(}f(x)-f(z)-\nabla f(z)\cdot(x-z)\big{)}.$
@@
(3.2) $\int_{\mathbb{R}^{d}}\big{|}|\xi|^{\alpha}\hat{g}(\xi)\big{|}^{2}\,\mathrm{d}\xi\leq CR^{d}.$
@@
Recall that, for $\alpha\in(1,2)$, $[\nabla f]_{\alpha-1}\lesssim\|D_{\alpha}\|_{*}$. Over the course of this proof, we will bound terms on the order of $[\nabla f]_{\alpha-1}^{2}R^{d}$, which is dominated by $\|D_{\alpha}f\|_{*}^{2}R^{d}$.
@@
$\int_{\mathbb{R}^{d}}\left(\hat{g}(\xi)|\xi|^{\alpha}\right)^{2}\,\mathrm{d}\xi\approx\int_{\mathbb{R}^{d}}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x.$
The display above uses standard facts about integral representation of the fractional Laplacian for $\alpha$ in the range $20<\alpha<$ 2. We bound expression on the right hand side above up to a harmless constant as follows
$\int_{|x-z|\geq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x+\int_{|x-z|<100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x=I+II.$
@@
(3.3) $I=\int_{|x-z|\geq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ =\int_{|x-z|\geq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{\phi\left(\frac{y-z}{R}\right)\big{(}f(y)-f(z)-\nabla f(z)\cdot(y-z)\big{)}}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ \lesssim_{[\nabla f]_{\alpha-1}^{2}}R^{2\alpha}\int_{|x-z|\geq 100R}\left(\int_{|y-z|\leq 4R}\frac{1}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ \lesssim R^{2d+2\alpha}\int_{|x-z|\geq 100R}\frac{1}{|x-z|^{2d+2\alpha}}\,\mathrm{d}x\approx R^{d}.$
@@
$g(x)-g(y)=\left(\phi\left(\frac{x-z}{R}\right)-\phi\left(\frac{y-z}{R}\right)\right)\left(f(x)-f(z)-\nabla f(z)\cdot(x-z)\right)\\ -\phi\left(\frac{y-z}{R}\right)\left(f(y)-f(x)-\nabla f(z)\cdot(y-x)\right).$
@@
$II\lesssim\int_{|x-z|<100R}\left(p.v.\int_{|x-y|\leq 10R}\frac{f(x)-f(z)-\nabla f(z)\cdot(x-z)}{|x-y|^{d+\alpha}}\left(\phi\left(\frac{x-z}{R}\right)-\phi\left(\frac{y-z}{R}\right)\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ +\int_{|x-z|<100R}\left(\int_{|x-y|\geq 10R}\frac{f(x)-f(z)-\nabla f(z)\cdot(x-z)}{|x-y|^{d+\alpha}}\left(\phi\left(\frac{x-z}{R}\right)-\phi\left(\frac{y-z}{R}\right)\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ +\int_{|x-z|<100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{f(y)-f(x)-\nabla f(z)\cdot(y-x)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ =:III+IV+V.$
@@
$|f(x)-f(z)-\nabla f(z)\cdot(x-z)|\lesssim_{[\nabla f]_{\alpha-1}}R^{\alpha},$
$\left|\phi\left(\frac{y-z}{R}\right)-\phi\left(\frac{x-z}{R}\right)-\nabla\left(\phi\left(\frac{x-z}{R}\right)\right)\cdot(y-x)\right|\lesssim\frac{|x-y|^{2}}{R^{2}}.$
@@
$\int_{|x-z|<100R}\bigg{(}p.v.\int_{|x-y|\leq 10R}\frac{f(x)-f(z)-\nabla f(z)\cdot(x-z)}{|x-y|^{d+\alpha}}\\ \cdot\left(\phi\left(\frac{x-z}{R}\right)+\nabla\left(\phi\left(\frac{x-z}{R}\right)\right)\cdot(y-x)-\phi\left(\frac{y-z}{R}\right)\right)\,\mathrm{d}y\bigg{)}^{2}\,\mathrm{d}x.$
@@
$[\nabla f]_{\alpha-1}^{2}\int_{|x-z|<100R}R^{2\alpha-4}\left(\int_{|x-y|\leq 10R}\frac{1}{|x-y|^{d+\alpha-2}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\lesssim[\nabla f]_{\alpha-1}^{2}R^{d}.$
@@
$\left|\phi\left(\frac{x-z}{R}\right)-\phi\left(\frac{y-z}{R}\right)\right|\lesssim\frac{|x-y|}{R}.$
@@
$IV\lesssim_{[\nabla f]_{\alpha-1}^{2}}\int_{|x-z|<100R}R^{2\alpha-2}\left(\int_{|x-y|\geq 10R}\frac{1}{|x-y|^{d+\alpha-1}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\lesssim R^{d}.$
@@
$V\lesssim\int_{|x-z|<100R}\left(\int_{|x-y|\geq 10R}\frac{(\nabla f(z)-\nabla f(x))\cdot(x-y)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ +\int_{|x-z|<100R}\left(p.v.\int_{|x-y|\leq 10R}\frac{(\nabla f(z)-\nabla f(x))\cdot(x-y)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ +\int_{|x-z|<100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{f(y)-f(x)-\nabla f(x)\cdot(y-x)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ =:VI+VII+VIII.$